# Variable Substitution and Bilinear Programming for Aligning Partially Overlapping Point Sets

### v1

@@ Abstract
In many applications, the demand arises for algorithms capable of aligning partially overlapping point sets while remaining invariant to the corresponding transformations. This research presents a method designed to meet such requirements through minimization of the objective function of the robust point matching (RPM) algorithm. First, we show that the RPM objective is a cubic polynomial. Then, through variable substitution, we transform the RPM objective to a quadratic function. Leveraging the convex envelope of bilinear monomials, we proceed to relax the resulting objective function, thus obtaining a lower bound problem that can be conveniently decomposed into distinct linear assignment and low-dimensional convex quadratic program components, both amenable to efficient optimization. Furthermore, a branch-and-bound (BnB) algorithm is devised, which solely branches over the transformation parameters, thereby boosting convergence rate. Empirical evaluations demonstrate better robustness of the proposed methodology against non-rigid deformation, positional noise, and outliers, particularly in scenarios where outliers remain distinct from inliers, when compared with prevailing state-of-the-art approaches.
@@ 1 Introduction
Aligning 2D/3D point sets is crucial across computer vision, robotics, and remote sensing applications. Apart from disturbances like non-rigid deformation and positional noise, this task becomes especially challenging when the two point sets only partial overlap and lack initial coarse alignment.
A practical approach to addressing the point set registration challenge involves minimizing the RPM objective [1]. This objective is characterized by a mixed linear assignment-least squares formulation:
$E(P,\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})\|^{2}$
Here, ${\mathbf{x}_{i}}$ and ${\mathbf{y}_{j}}$ represent two point sets intended for alignment, $p_{ij}\in{0,1}$ indicates point correspondence, and $\mathbf{T}$ denotes a spatial transformation parameterized by $\boldsymbol{\theta}$. Lian et al. demonstrated that in scenarios where one point set can be embedded within another, after spatial transformations are eliminated, the RPM objective can be reduced to a concave quadratic function of point correspondence, exhibiting a low-rank structure [2]. This simplification facilitates the implementation of efficient branch-and-bound (BnB) based global optimization strategies.
The algorithm introduced in [2] is extended to handle partial overlap in [3]. Here, the objective function takes the form of a non-quadratic concave function of point correspondence variable, exhibiting a low-rank structure. This structure enables application of efficient BnB optimization techniques. However, the algorithm’s branching space is a projected space of point correspondence, characterized by a dimensionality surpassing that of the transformation parameters. This results in a slower convergence rate for the method. Additionally, to ensure concavity of the objective function over the initial bounding region, the method requires the regularization of transformations, which in turn demands prior knowledge of transformation parameters. Consequently, the method lacks invariance to corresponding transformations, presenting a notable drawback.
@@
In response to this challenge, Lian and Zhang utilized the inherent constraints of rigid transformations to devise a novel objective function for the point correspondence variable. This function exhibits concavity across the entire solution space, eliminating the need for regularization of transformations [4]. Alternatively, Lian et al. employed the polyhedral annexation (PA) algorithm to minimize the RPM objective [5]. PA offers an advantage over BnB by operating exclusively within the concavity region of the objective, thus eliminating the need for transformation regularization. However, both approaches [4] and [5] still face the challenge inherited from [3] of branching within a high-dimensional projected space of point correspondence, resulting in sluggish convergence rates.
The methods mentioned above all share a common practice: they remove spatial transformations from the RPM objective, resulting in a function based solely on point correspondence variable. However, this simplification leads to a complex function, posing challenges such as high-dimensional branching spaces and intricate optimization strategies. An alternative approach, as demonstrated by [6], suggests retaining spatial transformations. In this method, trilinear and bilinear relaxation are employed to derive a lower bound function. This approach offers several advantages, including its ability to handle partial overlap, its invariance to corresponding transformations, and its utilization of a low-dimensional branching space.
Following the approach outlined in [6], we introduce a new Branch-and-Bound (BnB) algorithm tailored for optimizing the RPM objective function. This algorithm boasts the following features:
  • The algorithm proposed by [6] is complex because it necessitates reformulating the RPM objective function to ensure that one of the variables within the trilinear monomial includes the origin within its range. In contrast, as illustrated in Fig. 1, our method exclusively utilizes bilinear relaxation to establish the lower bound function, eliminating the necessity to fulfill this condition. This results in a streamlined algorithm compared to [6].
  • The optimization of the lower bound function can be broken down into two distinct tasks: linear assignment and low-dimensional convex quadratic optimization related to spatial transformation variables. Both of these tasks can be efficiently solved with global optimality assurances.
  • Our approach inherits the advantageous characteristics outlined in [6], including its ability in handling partial overlap, invariance to corresponding transformations, and utilization of a low-dimensional branching space.
The paper is structured as follows: We begin with a review of related work (Sec. 2). Following that, we introduce the theoretical framework necessary for developing our lower bound functions (Sec. 3). Next, we delve into our objective function (Sec. 4 and optimization strategy (5). Subsequently, we apply the proposed algorithm to 2D and 3D registration (Sec. 6 and Sec. 7). Finally, we present our experimental results (Sec. 8) and conclude the paper (Sec. 9).
@@ 2 Related work
A comprehensive survey of point set registration methods is available in [7]. Here, we focus solely on works closely related to ours.
@@ 2.1 Local methods
Simultaneous pose and correspondence. The widely known ICP method [8] iterates between recovering point correspondence and updating spatial transformation. Despite its efficiency, the discrete nature of point correspondence often traps the method in local minima. Zhang et al. enhanced the robustness of ICP by constraining the rotation angle at each iterative step [9]. Liang introduced RGB information to the registration process of ICP [10]. The RPM method [1] relaxes point correspondence to be fuzzily valued and employs deterministic annealing (DA) to gradually recover point correspondence. Instead of DA, Sofka et al. utilized the covariance matrix of the transformation parameters to determine the fuzziness of point correspondence, resulting in better robustness to missing or extraneous structures [11]. Ma et al. employed L2E to robustly estimate transformation for non-rigid registration [12]. Our method is related to the aforementioned methods by also modeling spatial transformation and correspondences. However, our objective function is globally minimized, which enhances our method’s robustness to disturbances.
Correspondence-free methods. The CPD method [13] frames point matching as fitting a Gaussian mixture model (GMM) representing one point set to another. GMMREG [14] utilizes two GMMs to represent two point sets and minimizes the $L_{2}$-norm distance between them. The SVGMs method represents point sets using sparse Gaussian components [15]. Efficiency in GMM-based methods is enhanced by employing filtering to solve the correspondence problem [16]. The density variation problem of point sets is addressed in [17] by modeling scene structure. Hierarchical Gaussian mixture representation improves speed and accuracy of registration [18].
@@ 2.2 Global methods
BnB based methods. The Branch-and-Bound (BnB) algorithm stands as a widely-used global optimization technique. Rooted in Lipschitz theory, BnB is applied to align 3D shapes [19], albeit with limitations regarding outliers or occlusions. By capitalizing on the structure of 3D rigid motions geometry, Go-ICP [20] globally optimizes the objective of ICP. GOGMA [21] achieves point set registration by aligning Gaussian Mixture Models (GMMs) constructed from the original point sets. Employing stereographic projection, the fast rotation search (FRS) method efficiently recovers 3D rotation between two point sets [22]. Bayesian nonparametric mixture and a novel approach of tessellating rotation space are introduced in [23], leading to an efficient alignment algorithm. Our BnB-based method distinguishes itself from the aforementioned methods by also incorporating point correspondences. This feature empowers our method to better handle non-rigid deformation.
Mismatch removal methods. An alternative avenue of research focuses on recovering spatial transformations given putative point correspondences. The fast global registration (FGR) method [24] optimizes a robust objective grounded in the duality between line processes and robust estimation. GORE [25] eliminates a large portion of outliers using geometric operations before invoking RANSAC. TEASER++ [26] employs a graph-theoretic approach to decouple rotation, scale, and translation estimation. It adopts a truncated least squares formulation for each subproblem, enabling effective mismatch removal.
2.3 Deep learning methods
Correspondence-free methods. The Lucas–Kanade algorithm aligns global features of point sets generated by a PointNet network in [27]. This approach is refined in [28] by employing a decoder to convert the generated global features back into point sets. Ensuring fidelity, Chamfer distances between the decoded point sets and the original ones are minimized.
Correspondence-based methods. The FCGF method [29] employs sparse high-dimensional convolutions to extract dense feature descriptors from point clouds. This method significantly outperforms many hand-crafted features and PointNet-based methods. Additionally, it surpasses a 3D convolution-based method [30]. The deep global registration method [31] utilizes FCGF feature descriptors for point cloud registration. DCP [32] utilizes DGCNN [33] and Transformer [34] to extract features for each point set, with SVD employed to recover rigid transformation. PRNet [35] extends DCP to handle partially overlapping point sets iteratively. IDAM [36] considers both geometric and distance features during its point matching process. RIENet [37] calculates feature-to-feature correspondences with neighborhood consensus. Finally, [38] performs partial point set registration by generating matching credibility on the global semantic level and learning geometric features on the local structural level.
@@ 3 Bilinear convex envelope linearization
The convex envelope of a bilinear monomial $xy$ subject to constraints $\underline{x}\leq x\leq\overline{x}$ and $\underline{y}\leq y\leq\overline{y}$ can be described as:
$\begin{cases}(xy)_{l1}\triangleq\underline{x}y+x\underline{y}-\underline{x}\underline{y}\leq xy,\\ (xy)_{l2}\triangleq\overline{x}y+x\overline{y}-\overline{x}\overline{y}\leq xy\end{cases}$ (1)
There are two ways of utilizing this formula when devising the lower bound problem: (1) Integration of the term $\max\{(xy)_{l1},(xy)_{l2}\}$ into the lower bound objective function. Nevertheless, this incorporation can lead to non-linearity of the resultant function. (2) Formulation of a subproblem in the format $\min\{\alpha|\alpha\geq(xy)_{l1},\alpha\geq(xy)_{l2}\}$ within the lower bound problem. However, this approach increases the complexity of constraints and raises concerns regarding the feasibility of the resulting problem for efficient optimization via combinatorial optimization algorithms.
Hence, in alignment with the methodology proposed in [6], we advocate for using the average of the convex envelope components, as delineated in (2):
$(xy)_{l}\triangleq\frac{1}{2}\sum_{i}(xy)_{li}\leq xy$ (2)
as the lower bound function for $xy$. This choice has the advantage that the lower bound function is linear, thereby facilitating efficient optimization by combinatorial optimization algorithms.
[6] also shows that for $(xy)_{l}$ to converge to $xy$, we only need to branch over one variable $x$ or $y$ that makes up the bilinear monomial $xy$. This is a crucial property and the good convergence of the proposed algorithm hings on this fact.
@@ 4 Problem formulation
Suppose there are two point sets $\mathscr{X}=\{\mathbf{x}_{i},1\leq i\leq n_{x}\}$ and $\mathscr{Y}=\{\mathbf{y}_{j},1\leq j\leq n_{y}\}$ in $\mathbb{R}^{n_{d}}$ to be aligned, where the column vectors $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ represent the coordinates of points $i$ and $j$, repectively. Based on [3], this problem can be modeled as a mixed linear assignment--least square problem:
@@
[…]
where
$\boldsymbol{\theta}$ : parameters of the spatial transformation $\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}$. $\mathbf{J}(\mathbf{x}_{i})$ : the Jacobian matrix at point $\mathbf{x}_{i}$. $\mathbf{I}_{d}$: the $d\times d$ identity matrix. $\mathbf{1}_{n_{x}}$: the $n_{x}$-dimensional vector of all ones. $\otimes$: the Kronecker product. $\mathbf{P}=\{p_{ij}\}$ : assignment matrix, with element $p_{ij}{=}1$ or 00 indicating that $\mathbf{x}_{i}$ matches $\mathbf{y}_{j}$ or not. $\mathbf{1}_{n_{x}}^{\top}\mathbf{P}\mathbf{1}_{n_{y}}{=}n_{p}$: the number of matches is a priori known to be a constant integer $n_{p}$. $\underline{\boldsymbol{\theta}}$ (resp. $\overline{\boldsymbol{\theta}}$) : the lower (resp. upper) bound of $\boldsymbol{\theta}$.
$\mathbf{y}\triangleq\begin{bmatrix}\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{n_{y}}^{\top}\end{bmatrix}^{\top}$: a vector. $\widetilde{\mathbf{y}}\triangleq\begin{bmatrix}\|\mathbf{y}_{1}\|_{2}^{2},\dots,\|\mathbf{y}_{n_{y}}\|_{2}^{2}\end{bmatrix}^{\top}$: a vector. $\mathbf{J}\triangleq\begin{bmatrix}\mathbf{J}^{\top}(\mathbf{x}_{1}),\dots,\mathbf{J}^{\top}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix.
For ease of derivation in the sequel, we need to convert matrix $\mathbf{P}$ into a vector $\mathbf{p}$. Accordingly, the objective function takes the following form [3]:
@@
where
$\mathbf{p}\triangleq\text{vec}(\mathbf{P})$: a vector. $\boldsymbol{\rho}\triangleq\mathbf{1}_{n_{x}}\otimes\widetilde{\mathbf{y}}$: a vector. $\mathbf{A}\triangleq(\mathbf{J}^{\top}\otimes\mathbf{y}^{\top})\mathbf{W}^{n_{x},n_{y}}_{n_{d}}$: a matrix. $\mathbf{B}\triangleq(\mathbf{J}_{2}^{\top}\otimes\mathbf{I}_{n_{\theta}})\mathbf{W}^{n_{x},1}_{n_{\theta}}(\mathbf{I}_{n_{x}}\otimes\mathbf{1}_{n_{y}}^{\top})$: a matrix. $\mathbf{J}_{2}\triangleq\begin{bmatrix}\mathbf{J}(\mathbf{x}_{1})^{\top}\mathbf{J}(\mathbf{x}_{1}),\ldots,\mathbf{J}(\mathbf{x}_{n_{x}})^{\top}\mathbf{J}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$: a matrix. $\mathbf{W}^{m,n}_{d}\triangleq\mathbf{I}_{m}\otimes\begin{bmatrix}\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{1})^{\top},\dots,\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{d})^{\top}\end{bmatrix}^{\top}$: a matrix.
$\text{vec}(\cdot)$: vectorization of a matrix by concatenating rows. $n_{\theta}$: the dimensionality of ${\boldsymbol{\theta}}$. $\mathbf{e}_{d}^{i}$: the $d$-dimensional column vector with only one nonzero unit element at the $i$-th position. $\text{mat}(\cdot)$: reconstructs a matrix from a vector, which is the inverse of the operator $\text{vec}(\cdot)$.
@@
where the nonzero elements of the constant matrix $\mathbf{C}$ correspond to the rows of $\mathbf{B}$ containing identical elements. The elements of the constant matrix $\mathbf{K}$ take values in $\{0,1\}$ and record the correspondences between the rows of $\mathbf{B}_{2}$ and those of $\mathbf{B}$.

### v2

# Decomposed Global Optimization for Robust Point Matching with Low-Dimensional Branching

@@ Abstract
Numerous applications require algorithms that can align partially overlapping point sets while maintaining invariance to geometric transformations (e.g., similarity, affine, rigid). This paper introduces a novel global optimization method for this task by minimizing the objective function of the Robust Point Matching (RPM) algorithm. We first reveal that the original RPM objective is a cubic polynomial. Through a concise variable substitution, we transform this objective into a quadratic function. By leveraging the convex envelope of bilinear monomials, we derive a tight lower bound for this quadratic function. This lower bound problem conveniently and efficiently decomposes into two parts: a standard linear assignment problem (solvable in polynomial time) and a low-dimensional convex quadratic program. Furthermore, we devise a specialized Branch-and-Bound (BnB) algorithm that branches exclusively on the transformation parameters, which significantly accelerates convergence by confining the search space. Experiments on 2D and 3D synthetic and real-world data demonstrate that our method, compared to state-of-the-art approaches, exhibits superior robustness to non-rigid deformations, positional noise, and outliers, particularly in scenarios where outliers are distinct from inliers.
@@ 1 Introduction
Point cloud registration, the process of aligning two point sets into a unified coordinate system, is a cornerstone of 3D computer vision, robotics, and augmented reality. While traditional methods like the Iterative Closest Point (ICP) algorithm are widely used, they’re notoriously sensitive to initial alignment and frequently get trapped in local minima, especially when dealing with significant outliers or partial data overlap. This fundamental limitation highlights the need for globally optimal algorithms that can reliably converge to the correct solution regardless of initialization.
The Branch-and-Bound (BnB) framework has emerged as a powerful tool for achieving such global guarantees. Many work applies BnB to solve the coupled problem of finding the optimal correspondence $\mathbf{P}$ and transformation $\mathbf{T}(\cdot|\boldsymbol{\theta})$, typically formulated as minimizing the following objective function:
$E(\mathbf{P},\boldsymbol{\theta})=\sum_{i,j}p_{ij}\|\mathbf{y}_{j}-\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})\|^{2},$ (1)
where $\mathbf{x}_{i}$} and $\mathbf{y}_{j}$} are points from the two sets, $p_{ij}\in\{0,1\}$ are binary variables indicating correspondence, and $\boldsymbol{\theta}$ parameterizes the spatial transformation. An Early attempt to solve this problem by branching over either $\mathbf{P}$ or $\boldsymbol{\theta}$ proved inefficient due to the immense search space and an inefficient bounding scheme. Subsequent research has largely split into two camps:
  1. 1.
    Correspondence-only Methods: These approaches eliminate the transformation variables $\boldsymbol{\theta}$, simplifying Eq. (1) into a problem that only involves correspondences. While this can reveal low-rank concave structures amenable to BnB optimization [1, 2], it comes at a cost. Initial methods were either restricted to subset alignment scenarios [1] or lacked transformation invariance [2]. While follow-up works [3, 4] successfully restored invariance, they could not overcome the core limitation of this approach: branching in the high-dimensional correspondence space. This large search space inherently leads to slow convergence, making these methods impractical for larger problems.
  2. 2.
    Transformation-space Methods: This more promising strategy retains the transformation parameters $\boldsymbol{\theta}$ and performs branching within the much smaller, low-dimensional transformation space. A representative example is the work by [5], which treats Eq. (1) as a cubic polynomial and uses trilinear relaxation to derive a lower bound for the BnB algorithm. This approach effectively handles partial overlap and maintains transformation invariance, benefiting from a significantly reduced search space.
1.1 Our Contributions
In this paper, we follow the second strategy of branching on the low-dimensional transformation space. However, we introduce a novel method to derive a simpler, yet more effective, lower bound that avoids a key complexity of prior work. The algorithm in [5] requires a complex reformulation of the objective function to ensure one of the variables in the trilinear monomial includes the origin within its range.
In contrast, as illustrated in Fig. 1, our method exclusively uses bilinear relaxation, which eliminates this requirement and results in a more straightforward and streamlined algorithm.
Our main contributions are summarized as follows:
  • We present a novel BnB algorithm for point cloud registration that utilizes a simpler, more robust lower-bounding scheme based on bilinear relaxation. This eliminates the need for complex objective function reformulations required by previous method [5], simplifying the overall algorithm.
  • Our lower bound optimization is efficient. It decomposes into two well-behaved subproblems: a classic linear assignment problem for correspondences and a low-dimensional convex quadratic program for the transformation. Both can be solved to global optimality very quickly.
  • Our resulting BnB algorithm inherits the key advantages of its predecessors, including robustness to partial overlap and invariance to transformations. Crucially, it leverages a low-dimensional branching space, enabling it to scale effectively to large problems while maintaining global optimality guarantees.
@@ 1.1 Our Contributions
The rest of the paper is organized as follows. Section 2 provides a review of related work. Section 3 introduces the theoretical framework for our lower bound. Section 4 details our objective function, and Section 5 describes the optimization strategy. We apply the algorithm to 2D and 3D registration in Sections 6 and 7, respectively. Finally, we present experimental results in Section 8 and conclude in Section 9.
@@ 2 Related Work
A comprehensive survey of point set registration is available in [6]. Here, we focus on works most closely related to our own, distinguishing between local and global optimization strategies.
@@ 2.1 Local Methods
These methods are efficient but do not guarantee global optimality, making them susceptible to local minima.
Simultaneous Pose and Correspondence. The venerable Iterative Closest Point (ICP) algorithm [7, 8] alternates between finding point correspondences and updating the spatial transformation. While efficient, its reliance on discrete assignments often leads it to get trapped in local minima. To enhance robustness, variations like that by Zhang et al. [9] constrain the rotation angle, while others incorporate additional information, such as RGB data [10].
The Robust Point Matching (RPM) method [11] addresses this by relaxing correspondences to have fuzzy values, using deterministic annealing (DA) to gradually refine them. Sofka et al. [12] used the covariance of transformation parameters to control this fuzziness, achieving better robustness to missing structures. For non-rigid registration, Ma et al. [13] employed L2E for robust transformation estimation.
Our method, like these approaches, jointly models spatial transformation and correspondence. However, we achieve greater robustness to various disturbances by guaranteeing global optimality, which is not possible with these local methods.
Correspondence-free Methods. These methods, such as Coherent Point Drift (CPD) [14], avoid explicit correspondence matching. CPD frames registration as fitting a Gaussian Mixture Model (GMM) from one point set to another. GMMREG [15] represents both point sets with GMMs and minimizes the L2 distance between them. Similarly, SVGMs [16] use sparse Gaussian components. GMM-based methods have been made more efficient with filtering techniques [17], while issues with point set density have been addressed by modeling scene structure [18] and using hierarchical representations [19]. These methods are computationally intensive and can still suffer from local minima, whereas our approach offers a global optimality guarantee.
@@ 2.2 Global Methods
These methods offer guarantees of finding the globally optimal solution, albeit at a higher computational cost.
BnB-based Methods. The Branch-and-Bound (BnB) algorithm is a powerful global optimization technique. An early application based on Lipschitz theory [20] was sensitive to outliers. More recent methods leverage the specific geometry of transformations. Go-ICP [21] globally optimizes the ICP objective by exploring the structure of 3D rigid motions. GOGMA [22] registers point sets by aligning their corresponding GMMs. Other works, such as the Fast Rotation Search (FRS) [23], efficiently find 3D rotations, while a Bayesian approach [24] uses a novel tessellation of the rotation space.
While these methods offer global guarantees, they often focus on specific problems (e.g., rigid motion) and do not handle correspondence explicitly, which can limit their applicability to non-rigid deformations. Our method differs by jointly and globally optimizing both transformation and correspondence, making it more versatile.
Mismatch Removal Methods. An alternative research direction focuses on finding the correct spatial transformation from an initial set of putative correspondences that may contain many errors. The Fast Global Registration (FGR) method [25] optimizes a robust objective based on line processes. GORE [26] prunes outliers using geometric constraints before applying RANSAC. TEASER++ [27] uses a graph-theoretic approach to decouple the estimation of rotation, scale, and translation, formulating each as a truncated least squares problem. TR-DE [28] also decomposes the 6-DOF transformation to reduce computational complexity within a two-stage BnB search. TEAR [29] uses a truncated entry-wise absolute residual function for outlier robustness.
These methods excel at handling high outlier rates but require an initial set of correspondences, which can be difficult to obtain robustly in highly cluttered or unstructured environments. Our method avoids this by simultaneously solving for both correspondence and transformation from scratch.
@@ 3 Bilinear Convex Envelope Linearization
Our approach to constructing a lower bound for the BnB algorithm relies on the convex envelope of a bilinear monomial $xy$. For variables $x$ and $y$ within a rectangular domain defined by $\underline{x}\leq x\leq\overline{x}$ and $\underline{y}\leq y\leq\overline{y}$, the convex envelope is the maximum of two linear functions:
$(xy)_{l}=\max\left\{\underline{x}y+x\underline{y}-\underline{x}\underline{y},\quad\overline{x}y+x\overline{y}-\overline{x}\overline{y}\right\}\leq xy.$ (2)
While this provides the tightest possible linear lower bound, using a ‘max‘ function within our objective would result in a non-linear problem, complicating the subproblem optimization.
To maintain linearity and ensure computational efficiency, we use a simpler, linearized lower bound. We use the average of the two linear components from the convex envelope:
$(xy)_{l}\triangleq\frac{1}{2}\left(\left(\underline{x}y+x\underline{y}-\underline{x}\underline{y}\right)+\left(\overline{x}y+x\overline{y}-\overline{x}\overline{y}\right)\right)\leq xy.$ (3)
This choice results in a strictly linear function, which is crucial for decomposing our lower bound subproblem into solvable components.
A key property of this average-based lower bound is its relationship to the branching process. As shown by [5], for the lower bound to converge to the true value of the monomial, it is sufficient to branch on only one of the variables—either $x$ or $y$. This property is fundamental to the efficiency of our algorithm, as it allows us to effectively reduce the search space by branching only on the transformation parameters $\boldsymbol{\theta}$ while leaving the correspondence variables $p_{ij}$ unbranched. This single-variable branching strategy is the cornerstone of our algorithm’s convergence and its ability to scale to larger problems.
@@ 4 Problem formulation
We aim to find a globally optimal alignment between two point sets, $\mathscr{X}=\{\mathbf{x}_{i}\}_{i=1}^{n_{x}}$ and $\mathscr{Y}=\{\mathbf{y}_{j}\}_{j=1}^{n_{y}}$ in $\mathbb{R}^{n_{d}}$. Following the mixed linear assignment-least squares framework of [2], the problem is formulated as minimizing the following objective function:
@@
$s.t.\ \mathbf{P}\mathbf{1}_{n_{y}}\leq\mathbf{1}_{n_{x}},\ \mathbf{1}_{n_{x}}^{\top}\mathbf{P}\leq\mathbf{1}_{n_{y}}^{\top},\ \mathbf{1}_{n_{x}}^{\top}\mathbf{P}\mathbf{1}_{n_{y}}=n_{p},\ \mathbf{P}\geq 0,\ \underline{\boldsymbol{\theta}}\leq\boldsymbol{\theta}\leq\overline{\boldsymbol{\theta}}$ (4c)
where $\mathbf{P}=\{p_{ij}\}$ is the binary assignment matrix, with $p_{ij}=1$ if point $\mathbf{x}_{i}$ matches $\mathbf{y}_{j}$, and 0 otherwise. The spatial transformation is parameterized by $\boldsymbol{\theta}$ such that $\mathbf{T}(\mathbf{x}_{i}|\boldsymbol{\theta})=\mathbf{J}(\mathbf{x}_{i})\boldsymbol{\theta}$, where $\mathbf{J}(\mathbf{x}_{i})$ is the Jacobian matrix at point $\mathbf{x}_{i}$. The constraint $\mathbf{1}_{n_{x}}^{\top}\mathbf{P} \mathbf{1}_{n_{y}}=n_{p}$ specifies that the number of matches is a known constant $n_{p}$, which models partial overlap. $\mathbf{I}_{d}$ is the $d\times d$ identity matrix. $\mathbf{1}_{n_{x}}$ is the $n_{x}$-dimensional vector of all ones. $\otimes$ is the Kronecker product. $\underline{\boldsymbol{\theta}}$ (resp. $\overline{\boldsymbol{\theta}}$) is the lower (resp. upper) bound of $\boldsymbol{\theta}$. Vectors $\mathbf{y}\triangleq\begin{bmatrix}\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{n_{y}}^{\top}\end{bmatrix}^{\top}$, $\widetilde{\mathbf{y}}\triangleq\begin{bmatrix}\|\mathbf{y}_{1}\|_{2}^{2},\dots,\|\mathbf{y}_{n_{y}}\|_{2}^{2}\end{bmatrix}^{\top}$ and matrix $\mathbf{J}\triangleq\begin{bmatrix}\mathbf{J}^{\top}(\mathbf{x}_{1}),\dots,\mathbf{J}^{\top}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$.
For the purpose of optimization, we reformulate the objective by vectorizing the matrices. By converting the assignment matrix $\mathbf{P}$ to a vector $\mathbf{p}=\text{vec}(\mathbf{P})$, the objective function can be written in a more compact form:
@@
Here, $\mathbf{A}\triangleq(\mathbf{J}^{\top}\otimes\mathbf{y}^{\top})\mathbf{W}^{n_{x},n_{y}}_{n_{d}}$, $\mathbf{B}\triangleq(\mathbf{J}_{2}^{\top}\otimes\mathbf{I}_{n_{\theta}})\mathbf{W}^{n_{x},1}_{n_{\theta}}(\mathbf{I}_{n_{x}}\otimes\mathbf{1}_{n_{y}}^{\top})$, and $\boldsymbol{\rho}\triangleq\mathbf{1}_{n_{x}}\otimes\widetilde{\mathbf{y}}$ are constant matrices and vectors derived from the point coordinates and Jacobian matrices, Matrices $\mathbf{J}_{2}\triangleq\begin{bmatrix}\mathbf{J}(\mathbf{x}_{1})^{\top}\mathbf{J}(\mathbf{x}_{1}),\ldots,\mathbf{J}(\mathbf{x}_{n_{x}})^{\top}\mathbf{J}(\mathbf{x}_{n_{x}})\end{bmatrix}^{\top}$ and $\mathbf{W}^{m,n}_{d}\triangleq\mathbf{I}_{m}\otimes\begin{bmatrix}\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{1})^{\top},\dots,\mathbf{I}_{n}\otimes(\mathbf{e}_{d}^{d})^{\top}\end{bmatrix}^{\top}$. $n_{\theta}$ denotes the dimensionality of $\boldsymbol{\theta}$}. $\mathbf{e}_{d}^{i}$ is the $d$-dimensional column vector with only one nonzero unit element at the $i$-th position. $\text{mat}(\cdot)$ reconstructs a matrix from a vector, which is the inverse of the operator $\text{vec}(\cdot)$.
@@
where the nonzero elements of the constant matrix $\mathbf{C}$ correspond to the rows of $\mathbf{B}$ containing identical elements. (Please refer to Sec. 6 and 7 for examples of $\mathbf{C}$). The elements of the constant matrix $\mathbf{K}$ take values in $\{0,1\}$ and record the correspondences between the rows of $\mathbf{B}_{2}$ and those of $\mathbf{B}$.